the reciprocal of 4 plus the reciprocal of 5 is the reciprocal of what number
Answer by richwmiller(17219)

Are we talking about multiplicative or addition reciprocal?
1/4+1/5=1/x
x=20/9
1/4+1/5=9/20
5/20+4/20=9/20
20/9 is the reciprocal of 9/20